Circular No. 6060

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

X-RAY NOVA IN SCORPIUS C. Bailyn, S. Jogee, and J. Orosz, Department of Astronomy, Yale

University, report: “We have continued to observe the optical nova reported on IAUC 6050. Revised astrometry yields a position of α = 16^{h}54^{m}00^{s}.13, δ = −39^{o}50′45′′.2 (equinox 2000.0), < 2′′ from the radio position reported by Hjellming (IAUC 6055) and well within the expected errors of the HST guide stars used as a basis of our astrometry. Between Aug. 10 and 16, the nova declined in magnitude from V = 14.5 to 15.0, while its color remained within 0.02 of V − I = +1.83. Further photometry of a possible eclipse: Aug. 17.018 UT, V = 14.74, V − I = +1.87; 17.068, 14.95, +1.94; 17.175, 14.90, +1.93; 17.221, 14.64, +1.83. We monitored the nova for a 5-hr period on each of the following four nights, during which time it was within 0.1 mag of V = 14.1 with a color between V − I = +1.80 and +1.85. Over a 4-hr span on Aug. 22, the nova was at V = 14.4. Our run at Cerro Tololo is now complete, and we strongly encourage other observers to monitor the further development of this interesting source.”

R. M. Hjellming, National Radio Astronomy Observatory, reports the results of 10-km Very Large Array radio observations of the field of GRO J1655-40: “Observations on Aug. 18.2 and 19.06 UT at frequencies of 1.4, 4.9, 8.4, 14.9, and 22.5 GHz show the total fluxes to be 5.5, 2.5, 1.6, 1.0, and 0.6 Jy, respectively, on the first day, and 3.9, 1.8, 1.3, 0.9, and 0.6 Jy on the second day. At all frequencies except 1.4 GHz, the radio source was resolved; on Aug. 18.2 its angular size was 0′′.2, and on Aug. 19.06, 0′′.3. The 22.5-GHz image on Aug. 19.06 shows an apparent double source with separation of 0′′.3 and a position angle of 42^{o}, with the southwest component twice the strength of the northeast component. If the expansion was linear, it began roughly on Aug. 16, and the apparent separation was increasing by 0′′.1 per day.”

COMET NAKAMURA-NISHIMURA-MACHHOLZ (1994m) Further total visual magnitude estimates: Aug. 5.90 UT, 8.5 (R. Haver,

Monte Serra, Italy, 15×80 binoculars); 8.95, 8.5 (A. Diepvens, Balen, Belgium, 0.15-m refractor); 11.24, 7.8 (C. E. Spratt, Victoria, BC, 0.20-m reflector); 16.02, 8.5 (Diepvens); 18.23, 8.1 (Spratt).

1994 August 24 Daniel W. E. Green